OpenStudy (anonymous):

which is an equation for the line with slope 1/3 through (-2,1)

OpenStudy (anonymous):

y=(1/3)x+ 5/3

OpenStudy (anonymous):

If you aren't looking for any specific type of equation, you could write it in point slope form, which is\[y-a = m(x-b)\] Where m is your slope. a is 1 and b is -2 because they relate as -2 = y and 1 = x